OpenStudy (anonymous):

Solve for x: −3|2x + 6| = −12

OpenStudy (yttrium):

divide the whole equation by (-3) first. it will become |2x+6| = 4 Then consier two cases: 1st: 2x+6 = 4 2nd: -(2x+6) = 4 Solve for x and you'll get the answer.

OpenStudy (anonymous):

ok so this is what I got far -11 and -3 ??

OpenStudy (anonymous):

oh never mind that's wrong

OpenStudy (anonymous):

I made a mistake

OpenStudy (anonymous):

so -2 and -10

OpenStudy (anonymous):

?

OpenStudy (nincompoop):

what do you know about absolute value?

OpenStudy (anonymous):

That its the distance from zero ?

OpenStudy (yttrium):

-10 is correct. But are you sure with -2??

OpenStudy (anonymous):

positive 2 ?

OpenStudy (yttrium):

Nope. Try to solve it again. :)

OpenStudy (anonymous):

ok

OpenStudy (haseeb96):

may i solve your question completely?

OpenStudy (anonymous):

I got -1 & yes

OpenStudy (anonymous):